OpenStudy (anonymous):

Solve by factoring: 6x2 + x = 5

OpenStudy (anonymous):

6x^2 + x = 5 or 6x^2 + x - 5 = 0 or 6x^2 + 6x - 5x - 5 = 0 or 6x(x + 1) - 5(x + 1) = 0 or (6x - 5)(x + 1) = 0 or x = 5/6, -1

OpenStudy (anonymous):

6x^2+x-5=0 x=